How much electrical energy is required to generate 1.000L of H, at 25°C and 1.00 atm pressure by the electrolysis of water?


1
Expert's answer
2021-01-22T06:19:09-0500

Redox reactions in water electrolysis :

Anode: 2H2O (l) = O2 (g) + 4H+ (aq) + 4e-

Cathode: 4H2O (l) + 4e- = 2H2 (g) + 4OH- (aq)

Reaction: 2H2O = 2H2 (g) + O2 (g)

mol H2:

n = PV/RT

n = (1.00 atm x 1.000 L)/(0.082 atm•L/mol•K x 298 K)

n = 0.041 mol

From reaction at cathode : mol ratio of electron and H2 = 4 : 2, so mol electrons :

= 4/2 x mol H2 = 2 x 0.041 mol = 0.082 mol

1 Faraday is the amount of electricity that is passed in the electrolysis cell to obtain 1 mole of electrons.

1 Faraday = 1 mole of electrons = 96500 Coulombs

So the energy required :

0.082 Faraday